OpenStudy (anonymous):

what is the x intercept of 3x+2y-3=0

OpenStudy (anonymous):

x=1. Use y=0.

OpenStudy (anonymous):

Let y = 0 and substitute.

OpenStudy (anonymous):

Yeah but what do i do with -3

OpenStudy (anonymous):

3x + 2(0) - 3 = 0 Now solve for x by adding three to both sides 3x - 3 + 3 = 0 + 3 3x = 3 and divide both sides by 3 x = 1

OpenStudy (anonymous):

okkk thanks